### v3

@@ 5.1.2 Various Approaches and Tools are Used to Mine Architectural Information from Diverse Sources
Regarding the proposed and employed approaches, we found that the research on mining architectural information mainly focuses on automatically mining architectural information from various sources (i.e., 60%, 57 out of 95 approaches). On the other hand, the results show that 56 tools have been proposed and utilized for mining architectural information, of which 30.4% (17 out of 56) are general tools and 69.6% (39 out of 56) are dedicated tools. The general tools are usually employed to support the dedicated tools in mining architectural information. Architectural information classification-based approaches 52.6% (50 out of 95) and dedicated tools that support architectural information classification (35.9%, 14 out of 39 tools) are the most frequently used approaches and tools in mining architectural information. One potential reason could be that, classification based approaches and tools normally require an existing classification, which is the common situation in the scenarios of mining architectural information, like the ISO 42010:2011 standard (ISO/IEC/IEEE, 2011) for the classification of architecture elements and the architectural decision ontology (Kruchten, 2004) for the classification of architectural decisions. Moreover, we observed that many dedicated tools (see the Automatic Tools and Semi-automatic tools sheets in the dataset (de Dieu et al., 2023)) used to mine architectural information are implemented to support architectural information mining approaches (see the Automatic Approaches, Semi-automatic Approaches, and Manual Approach sheets in the dataset (de Dieu et al., 2023)). However, for the proposed architectural information mining approaches (i.e., 95 approaches), still more than half of the approaches (i.e., 56 approaches) are not supported by tools.
The relationships between the approaches and tools (results of RQ4) and sources used (results of RQ2) for mining architectural information: We show the relationships between the tasks (e.g., classification tasks) in mining architectural information, architectural information mining approaches and tools, sources used for mining architectural information, and mined architectural information in the Automatic Approaches, Semi-automatic Approaches, Manual Approach, Automatic Tools, and Semi-automatic Tools sheets in the dataset (de Dieu et al., 2023)). These sheets can provide suggestions to researchers and practitioners about which categories of architectural information (e.g., architectural change, design relationship) can be mined by which architectural information mining approaches and tools from which sources (e.g., App Stores, developer mailing lists) of software repositories.
@@ 5.2 Implications for Researchers and Practitioners
For researchers, we believe that this SMS contributes by synthesizing and advancing the existing knowledge on mining architectural information in software development through following a reproducible research process in line with the guidelines defined for literature mapping (see Section 3). Moreover, our SMS highlights the potential for further research in the field of mining architectural information, suggesting that there are still unaddressed questions, such as why certain types of architectural information are less or not mined yet and what are the challenges to be handled. These challenges represent open opportunities for future research on mining architectural information.
For practitioners, this SMS provides the information to know and adopt the state-of-the-art research results, such as what approaches and tools can be used for mining what architectural information from what sources to support architecting process and help to utilize the types of architectural information that have been mined to support the development. Inspired by the results we obtained in this SMS, we list a set of suggestions that we concluded from the literature on mining architectural information in software development.
5.2.1 Implications of RQ1 Results
The results of RQ1 show that, although 7 categories and 29 subcategories of architectural information have been mined to support various architecting activities (results of RQ3), certain types of architectural information, such as design relationship, architectural change (e.g., adaptive change and preventive change), and architectural technical debt (e.g., architectural compliance issue), are rarely mined. For example, there has been less attention to information related to design relationship (see Table 5). An architecture of a system is designed with the aid of architectural elements (e.g., tactics and components). Nevertheless, there are complex interdependency relationships between these elements, and many of the tasks faced by developers are to efficiently and effectively use these elements together during the application design to meet both functional and non-functional goals (Karthik and Medvidovic, 2019; Bi et al., 2018). For example, designing with a certain architecture pattern to address specific quality requirements cannot be considered in isolation, but there might be other key architectural elements to be considered, such as design contexts (Bedjeti et al., 2017). Therefore, we encourage the software architecture community to further investigate (e.g., through conducting industrial studies) the causes and reasons (e.g., difficulties in developing dedicated approaches) that limit researchers to mine the rarely mined architectural information in order to support the development.
Suggestion 1
5.2.2 Implications of RQ2 Results
The results of RQ2 show that some sources are rarely used in mining architectural information, such as Code search engines (e.g., Searchcode), Search engines (e.g., Google), App Stores (e.g., Google Play Store), and Technical blogs and tutorials. Some potential reasons could be, for example, ambiguity, redundancy, and incompleteness in architectural element descriptions in certain sources (as reported in Section 4.6). As a future direction, we suggest that the software architecture community could further explore the reasons (e.g., unsuitability for mining architectural information, usefulness of architectural information) that restrain researchers and practitioners from mining those sources. For example, the evidence in the literature on mining architectural information reveals that technical blogs and tutorials are a common source of architecture knowledge on the Web (Soliman et al., 2021c). However, technical blogs and tutorials are currently not empirically researched for their architecture knowledge, and thus further investigation on architectural knowledge shared in technical blogs and tutorials is needed. Researchers might study the reasons, such as the quality of architectural knowledge shared in technical blogs and tutorials, that cause technical blogs and tutorials to be rarely mined.
Suggestion 2
Architectural information is dispersed in various sources of repositories, issue tracking systems (e.g., Jira) (Bhat et al., 2017), and technical blogs and tutorials (Soliman et al., 2021c). There is no single source of architectural information that contains specific architectural information concepts, and certain architectural information concepts come more often in certain sources (Soliman et al., 2021c)). For instance, developer communities (as a prominent example of architectural information source) contain predominantly general architectural information concepts (Tang et al., 2010), such as architectural solutions and rationale (e.g., the benefits and drawbacks of architectural solutions) (Soliman et al., 2017). In contrast, issue tracking systems contain mainly architectural issues (e.g., architectural debt) and design decisions taken to address those issues (Bhat et al., 2017). Moreover, architectural solutions and benefits & drawbacks of these solutions are less present within source code repositories, such as GitHub (Soliman et al., 2021c). Therefore, the selection of a source is determined by the type of architectural information to be mined as well as the purpose of mining architectural information. Based on the analysis conducted for RQ2 results, in Table 7, we suggest the most prominent sources and prevalent types of architectural information these sources contain. The findings presented in Table 7 can guide practitioners to the suitable sources for mining certain types of architectural information efficiently. For instance, to mine architectural solutions and rationale, such as benefits and drawbacks of architectural solutions, practitioners could give a higher priority to Q&A sites, Wikis, and technical blogs and tutorials. In addition, to mine information related to architectural issues and design decisions taken for addressing such issues may select issue tracking systems, such as Jira (see Table 7).
Table 7: Prominent sources for mining architectural information
Source Dataset type Prevalent types of architectural information in the source
Version control systems (e.g., GitHub)
Code-based dataset, Text-based dataset
Architectural description (e.g., architectural view), Architectural change (e.g., perfective change)
Online Q&A sites (e.g., Stack Overflow)
Text-based dataset
Architectural solution (e.g., architectural pattern), Architectural rationale (e.g., benefits and drawbacks of architectural solutions)
Wiki
Text-based dataset
Architectural solution (e.g., architectural pattern), Architectural rationale (e.g., benefits and drawbacks of architectural solution), System requirement (i.e., functional and non-functional requirements)
Issue tracking systems (e.g., Jira)
Text-based dataset, Code-based dataset
Architectural issue (e.g., architectural debt), Architectural decision (e.g., structural and behavioral decisions)
Online books, research articles, and white papers
Text-based dataset
Architectural description (e.g., architectural model, architectural view), Architectural solution (e.g., architectural pattern and tactic), Design relationship
Technical blogs and tutorials
Text-based dataset
Architectural solution (e.g., architectural pattern and tactic), architectural rationale (e.g., benefits and drawbacks of architectural solutions)
Suggestion 3
We suggest that practitioners should pay much attention on those prominent sources (see Table 7) when selecting sources for mining architectural information. For example, to collect architectural debt and design decisions taken for addressing such debt, practitioners could give a higher priority to issue tracking systems, such as Jira.
5.2.3 Implications of RQ3 Results
Although 7 categories and 29 subcategories of architectural information have been mined (results of RQ1) to solve various problems in different architecting activities (results of RQ3), we observed that some types of architectural information are not yet (or rarely) mined for solving the problems in certain architecting activities (e.g., AME and AIm). In the following paragraphs, we illustrate some of the unresolved problems in certain architecting activities.
Unresolved problems in AME - Undetected architectural technical debt: ATD is a critical type of architectural information for AME (Li et al., 2014). ATD compromises system-wide quality attributes, particularly maintainability and evolvability (Li et al., 2014), and ATD needs to be identified and well managed since it is harmful to the system’s long-term health (Li et al., 2015). We found that most studied architectural smells (one form of architectural technical debt) include Unstable Interface, Cyclic Dependency, Hub-like Dependency, Modularity Violations, Package Cycles, Unhealthy Inheritance Hierarchy, Chatty Service [S29][S71][S72] [S97][S96]. However, there is a variety of undetected architectural smells that are defined in the literature (Azadi et al., 2019), such as Scattered Functionality, God Component, Multipath Hierarchy, Multiple Architecture Violations, Unutilized Abstraction, Specification-Implementation Violation, and Dense Structure. Thus, the studies on detecting and mining these undetected architectural smells for facilitating AME activity would be a timely contribution to architecture research and practice. Refactoring detected architectural technical debt: We noticed that the majority of the studies (e.g., [S65][S100][S104]) in AME activity researched on the detection of architectural technical debt without proposing and recommending how to refactor the identified architectural debt. Therefore, approaches and tools for refactoring the detected architectural smells would be valuable for the developers.
Suggestion 4
Unresolved problems in AIm - Lack of broader focus on implementation of architectural elements: We found that a few studies researched on the implementation of certain architectural elements, such as architectural tactic implementations (e.g., [S13][S49]) in order to assist AIm activity. In most of the cases, these studies analyzed code bases for mining and reusing code snippets that implement architectural tactics (e.g., [S13]) during AIm. Moreover, we also noticed that there is a lack of studies on the implementation of other architectural elements, including architectural pattern and framework implementations. Therefore, future investigation into mining (e.g., via searching) and reusing the implementation of those architectural elements can speed up the development. For example, such studies can develop approaches and tools for searching and reusing code snippets that implement architectural patterns. Architecturally-relevant code summarization: This is another unresolved and interesting research direction in AIm activity. Architecturally-relevant code summarization is a task of writing brief natural language descriptions of architecturally-relevant code. The idea is that a short description could allow a programmer to understand what a section of architecturally-relevant code does and purpose of that code in the overall program, without requiring the programmer to read the code itself. However, we observed that there is no investigation on this research direction and such investigation can help developers effectively implement architectural elements during software development. Generation of architecturally-relevant code comments: We also observed that there are no studies on generation and summarization of architecturally-relevant code comments. Architecturally-relevant code comments are crucial to program comprehension and could speed up architecture implementation. If architecturally-relevant code generation is to automatically synthesize syntactically correct architecturally-relevant code based on developers’ intentions, then architecturally-relevant code comment generation is to produce one or several sentences to describe the semantic of architecturally-relevant code fragments. Therefore, it is important to conduct investigations in this direction to develop reliable approaches and tools (e.g., plugin tools embedded in IDE) in order to assist AIm activity.
Suggestion 5
Unresolved problems in AA - ASR traceability and validation in the code: As stated in Section 4.4, the main task in AA activity is to examine and identify ASRs that an architecture of a system need to satisfy. We observed that most of the studies in AA have analyzed and mined requirements (e.g., ASRs) from textual artifacts, such as requirement documentations and architectural documentations. However, requirements found in documentation can rarely be traced back to source code. This implies that documented requirements provide only limited support for keeping developers informed about the underlying architectural decisions during the development. Therefore, we believe that mining approaches and tools on traceability and validation of ASRs in code applied to a significant set of systems could help developers get insights into how architectural concerns are addressed in code.
Suggestion 6
5.2.4 Implications of RQ4 Results
Although architectural information mining approaches and tools (the results of RQ4) are useful in development (e.g., ArchEngine [S13]), as reported in Section 4.5, there is a demand to evaluate these approaches and tool in industry practice. We encourage practitioners to explore the convincing benefits and limitations by applying architectural information mining approaches and tools in industrial practice. Besides, researchers can evaluate the approaches and tools reported and explore their scope, characteristics, and metrics to provide a solid foundation for designing dedicated tools. Such evaluation could enhance the desire from practitioners and adopt the proposed approaches and tools for their specific needs in software development.
Suggestion 7
From the analysis of the results of RQ4 in Section 5.1.2, we can see that, for the proposed architectural information mining approaches (i.e., 95 approaches), still more than half of the approaches (i.e., 56 approaches) are not supported by tools. Thus, researchers and practitioners should examine the proposed architectural information mining approaches and develop architectural information mining tools to better support those approaches. For instance, in [S78], Kamal and Avgeriou proposed a manual approach for mining architectural pattern-architectural pattern relationships from books, research articles, and white papers, in order to assist architects in reusing architectural patterns during architecting process. When this approach can be (semi)-automated by a dedicated tool in architecting process, the efficiency of architecting activities, such as architectural reuse, maintenance, and evolution, will be significantly improved.
Suggestion 8
@@ 5.2.4 Implications of RQ4 Results
Despite numerous approaches and tools proposed for mining architectural information (see Section 4.5), there is no clear answer as to which architecting activities the proposed mining approaches and tools can be best applied and how to leverage the suitable approaches and tools for these activities. Answering these questions would be beneficial for both practitioners and researchers to make informed decisions to solve problems or conduct research using architectural information mining approaches and tools. Thus, based on the analysis conducted in this study, in Table 8 and Table 9, we provide the most promising approaches and tools for mining architectural information. Specifically, the information presented in Table 8 and Table 9 can guide practitioners know and adopt promising approaches and/or tools for mining certain types of architectural information from particular sources in order to support specific architecting activities. It is worth noting that we referred to the following factors (e.g., evidence levels, see Section 4.5) to identify the most promising approaches and tools provided in Table 8 and Table 9. Adaptation in industrial practice (Level 6): Some tools are currently being utilized in industry settings, such as DV8, ADeX, Arcan. Evidence obtained from industrial studies (Level 5): An approach or tool has been evaluated using case studies in industry with practitioners. For example, Juergen et al. [S53] proposed an approach named Continuous Architectural Knowledge Integration (CAKI) that combines the continuous integration of internal and external architectural knowledge sources with enhanced semantic searching, reasoning, and personalization capabilities dedicated to large organizations. The approach was evaluated in an industrial setting with practitioners from Siemens. Extensive evaluation on non-trivial industrial software systems: Some approaches and tools have been validated on a bunch of non-trivial industrial software projects. For example, in [S13], Mujhid et al. proposed a tool (i.e., ArchEngine) for searching and reusing architecturally significant code. The tool was evaluated on a large scare repository that contains 116,609 projects extracted from GitHub, Google Code, SourceForge, Apache, and other software repositories. In addition, the authors also evaluated the tool through a case study with 21 graduate students who have experience level of junior software developers and the evaluation results showed that ArchEngine is more effective than other search engines (e.g., Krugle and Koders).
@@
ARCADE
Architectural description (e.g., Architectural view)
Version control systems
Architectural recovery, Architectural maintenance and evolution
[S46]
CAKI
Architectural solution (e.g., Architectural pattern, Architectural tactic), Architectural decision (e.g., Technology decision)
Mixed sources (VCS, Wiki, Q&A sites, Online books, research articles, etc.)
Architecture Recovery
[S53]
Tactic Detector(DT)
Architectural solution (e.g., Architectural tactic)
VCS, Code search engines
Architecture reuse, Architecture implementation
[S57]
ArchDRH
Design relationship (e.g., Component-Component relationship)
VCS
Architecture recovery, Architecture maintenance and evolution
[S57]
Architecture Anti-pattern Detector
Architectural technical debt (e.g., Architectural anti-pattern)
VCS, Issue tracking systems
Architecture maintenance and evolution
[S64]
MaRK
System requirement (e.g., Functional requirement)
Wiki
Architecture implementation
[S45]
DARCY
Architectural technical debt (e.g., Architectural compliance issue), Architectural description (e.g., Architectural model)
VCS
Architecture conformance checking, Architecture understanding, Architecture maintenance
[S30]
ArCode
Architectural solution (e.g., Architectural tactic)
VCS
Architecture implementation
[49]
RecovAr
Architectural decision (e.g., Behavioral decision, Structural decision)
VCS
Architectural recovery
[S47]
@@
Tool Architectural information Source and dataset Supported architecting activities Studies
Arcan
Architectural technical debt (e.g., Architectural smell)
Issue tracking systems (Code-based dataset)
Architectural maintenance and evolution
[S71]
DV8
Architectural technical debt (e.g., Architectural smell)
VCS (Code-based dataset)
Architectural maintenance and evolution
[S94] [S96]
ADeX
Architectural decision (e.g., Structural, Behavioral, Ban or non-existence decisions)
Issue tacking systems (Text-based dataset)
Architecture understanding
[S19]
Hotspot detector
Architectural technical debt (e.g., Architectural smell)
VCS (Code-based dataset)
Architectural maintenance and evolution
[S93]
ArchEngine
Architectural solution (e.g., Architectural tactic)
VCS (Code-based dataset)
Architecture implementation
[S13]
Titan
Architectural description (Architectural view, Architectural model), Architectural technical debt (Architectural anti-pattern)
Issue tracking system (Code-based dataset)
Architecture maintenance and evolution
[S37]
Archie
Architectural solution (e.g., Architectural tactic)
Code search engines
Architecture reuse, Architecture implementation
[S57]
Suggestion 9
We suggest that practitioners could give high priorities to promising approaches and tools presented in Table 8 and Table 9 when selecting approaches and tools for mining architectural information from software repositories.
From the results of RQ4 (see Section 4.5), we observed that most primary studies motivated their mining approaches to reduce the manual effort in mining architectural information. Such rationale can be reasonable in the context of large-scale collections of software artifacts. Furthermore, manually mining relevant architectural information from large volume of software artifacts is not cost-effective. However, in our recent industrial survey on how developers search for architectural information (de Dieu et al., 2022), we found that none of the participants (practitioners) used the existing approaches and tools (proposed in the literature) to search and mine architectural information from software repositories. Therefore, it seems legitimate to study the potential impact of the architectural information mining research on the industry; and to what extent the architectural information mining approaches and tools would be useful in industrial settings. Such a study could address this problem from multiple perspectives, for example, what small, medium, and large software development organizations are interested in architectural information mining approaches and tools? Who in the organization would use the approaches or tools? Is mining architectural information manually ‘the real pain’ of the practitioners? if so, how does ‘the pain’ manifest itself? Are any tasks obstructed? Is the problem generating additional costs? Answering the questions could help to understand who are the actual beneficiaries of the architectural information mining research; and what is the size of that market. Not only it would help to scope and justify future research directions, but it would also provide insights into commercializing this research.
Suggestion 10
5.2.5 Implications of RQ5 Results

### v4

@@ 5.1.2 Various Approaches and Tools are Used to Mine Architectural Information from Diverse Sources
Regarding the proposed and employed approaches, we found that the research on mining architectural information mainly focuses on automatically mining architectural information from various sources (i.e., 60%, 57 out of 95 approaches). On the other hand, the results show that 56 tools have been proposed and utilized for mining architectural information, of which 30.4% (17 out of 56) are general tools and 69.6% (39 out of 56) are dedicated tools. The general tools are usually employed to support the dedicated tools in mining architectural information. Architectural information classification-based approaches 52.6% (50 out of 95) and dedicated tools that support architectural information classification (35.9%, 14 out of 39 tools) are the most frequently used approaches and tools in mining architectural information. One potential reason could be that, classification based approaches and tools normally require an existing classification, which is the common situation in the scenarios of mining architectural information, like the ISO 42010:2011 standard (ISO/IEC/IEEE, 2011) for the classification of architecture elements and the architectural decision ontology (Kruchten, 2004) for the classification of architectural decisions. Moreover, we observed that many dedicated tools (see the Automatic Tools and Semi-automatic tools sheets in the Supplementary Material (de Dieu et al., 2024)) used to mine architectural information are implemented to support architectural information mining approaches (see the Automatic Approaches, Semi-automatic Approaches, and Manual Approach sheets in the Supplementary Material (de Dieu et al., 2024)). However, for the proposed architectural information mining approaches (i.e., 95 approaches), still more than half of the approaches (i.e., 56 approaches) are not supported by tools.
The relationships between the approaches and tools (results of RQ4) and sources used (results of RQ2) for mining architectural information: We show the relationships between the tasks (e.g., classification tasks) in mining architectural information, architectural information mining approaches and tools, sources used for mining architectural information, and mined architectural information in the Automatic Approaches, Semi-automatic Approaches, Manual Approach, Automatic Tools, and Semi-automatic Tools sheets in the Supplementary Material (de Dieu et al., 2024)). These sheets can provide suggestions to researchers and practitioners about which categories of architectural information (e.g., architectural change, design relationship) can be mined by which architectural information mining approaches and tools from which sources (e.g., App Stores, developer mailing lists) of software repositories.
@@ 5.2 Implications for Researchers and Practitioners
5.2.1 Addressing unresolved problems in architecting activities
Although various types of architectural information have been mined (results of RQ1) to solve the problems in different architecting activities (results of RQ3), there are still unresolved problems in certain architecting activities, including architectural implementation (AIm) and maintenance & evolution (AME). For example, we found that a few studies focused on mining and recommending the implementation of certain architectural elements, such as architectural tactic implementations (e.g., [S13][S49]) in order to assist AIm activity. Furthermore, we also noticed that there is a lack of studies on the implementation of other architectural elements, including architectural pattern and framework implementations.
[…]
5.2.2 Promising approaches and tools for mining architectural information
@@
Despite numerous approaches and tools proposed for mining architectural information (see Section 4.5), there is no clear answer as to which architecting activities the proposed mining approaches and tools can be best applied and how to leverage the suitable approaches and tools for these activities. Answering these questions would be beneficial for both practitioners and researchers to make informed decisions to solve problems or conduct research using architectural information mining approaches and tools. Thus, based on the analysis conducted in this study, in Table LABEL:SuggestingApproachesToPractitioners and Table LABEL:SuggestingToolsToPractitioners, we provide the most promising approaches and tools for mining architectural information. Specifically, the information presented in Table LABEL:SuggestingApproachesToPractitioners and Table LABEL:SuggestingToolsToPractitioners can guide practitioners know and adopt promising approaches and/or tools for mining certain types of architectural information from particular sources in order to support specific architecting activities. It is worth noting that we referred to the following factors (e.g., evidence levels, see Section 4.5) to identify the most promising approaches and tools provided in Table LABEL:SuggestingApproachesToPractitioners and Table LABEL:SuggestingToolsToPractitioners. Adaptation in industrial practice (Level 6): Some tools are currently being utilized in industry settings, such as DV8. Evidence obtained from industrial studies (Level 5): An approach or tool has been evaluated using case studies in industry with practitioners. For example, Juergen et al. [S53] proposed an approach named Continuous Architectural Knowledge Integration (CAKI) that combines the continuous integration of internal and external architectural knowledge sources with enhanced semantic searching, reasoning, and personalization capabilities dedicated to large organizations. The approach was evaluated in an industrial setting with practitioners from Siemens. Extensive evaluation on non-trivial industrial software systems: Some approaches and tools have been validated on a bunch of non-trivial industrial software projects. For example, in [S30], Ghorbani et al. proposed an approach named DARCY for detecting architectural inconsistencies in Java software systems. The approach was evaluated on 38 open-source projects.
[…]
@@
ARCADE Architectural description (e.g., Architectural view) VCS Architectural recovery, Architectural maintenance and evolution [S46]
CAKI Architectural solution (e.g., Architectural pattern, Architectural tactic), Architectural decision (e.g., Technology decision) Mixed sources (VCS, Wiki, Q&A sites, Online books, research articles, etc.) Architecture Recovery [S53]
Architecture Anti-pattern Detector Architectural technical debt (e.g., Architectural anti-pattern) VCS, Issue tracking systems Architecture maintenance and evolution [S64]
ArchDRH Design relationship (e.g., Component-Component relationship) VCS Architecture recovery, Architecture maintenance and evolution [S101]
Tactic Detector(TD) Architectural solution (e.g., Architectural tactic) VCS, Code search engines Architecture reuse, Architecture implementation [S57]
DARCY Architectural technical debt (e.g., Architectural compliance issue), Architectural description (e.g., Architectural model) VCS Architecture conformance checking, Architecture understanding, Architecture maintenance [S30]
@@
Tool Architectural information Source and dataset Supported architecting activities URL link Studies
Arcan Architectural technical debt (e.g., Architectural smell) Issue tracking systems (Code-based dataset) Architectural maintenance and evolution https://essere.disco.unimib.it/wiki/arcan/ [S88]
DV8 Architectural technical debt (e.g., Architectural smell) VCS (Code-based dataset) Architectural maintenance and evolution https://archdia.com/ [S94]
5.2.3 Developing tools to better support the approaches
From the analysis of the RQ4 results in Section 5.1.2, we can see that, for the proposed architectural information mining approaches (i.e., 95 approaches), still more than half of the approaches (i.e., 56 approaches) are not supported by tools. For instance, in [S78], Kamal and Avgeriou proposed a manual approach for mining architectural pattern-architectural pattern relationships from books, research articles, and white papers, in order to assist architects in reusing architectural patterns during the architecting process. (Semi-)automating this approach with a dedicated tool can significantly improve the efficiency of architecting activities, such as architectural reuse (AReu), maintenance and evolution (AME).
[…]
5.2.4 Verifying the industrial needs for mining architectural information
From the results of RQ4 (see Section 4.5), we observed that most primary studies motivated their mining approaches to reduce the manual effort in mining architectural information. Such rationale can be reasonable in the context of large-scale collections of software artifacts. Furthermore, manually mining relevant architectural information from a large volume of software artifacts is not cost-effective. However, we have recorded no research studying what features the approaches or tools should possess nor what goals they should satisfy when mining architectural information from practitioners’ perspectives. Thus, the current research mainly satisfies requirements coming from scholars rather than practitioners. This indicates that the current research is intuition-based rather than practitioner-oriented. These studies applied different types of approaches and tools to mine architectural information from software repositories without explicitly examining the practitioners’ needs, and it is not clear to what extent the approaches and tools satisfy the real practitioners’ needs. Although some existing studies provided evidence levels (e.g., evidence obtained from industrial studies [S53][S45][S55][S64]) of how practitioners found certain architectural information mining approaches and tools useful, more systematic research is necessary in such directions to understand practitioners’ needs.
[…]
5.2.5 Open source benchmark architectural datasets for architecture research